Patrick invested %5200 for 21 years. At the investment's maturity, it's value was $13280.80 What was the annual simple interest rate?
A) 155.4% B) 7.4% C) 8.2% D) 2.6%

Answer by jorel1380(3719)

13280.8-5200=8080.80
8080.80/5200=1.554
1.554/21=.074, or 7.4% annual interest
☺☺☺☺